OpenStudy (amtran_bus):

Solve 18 + |2x − 9| = 17

OpenStudy (amtran_bus):

Is it no solution, false?

OpenStudy (arabpride):

18 + |2x − 9| = 17 -18 -18 |2x + 9| = -1 -9 -9 |2x| = -10 /2 /2 |x| = -10 |x| = |-10| x = 10

OpenStudy (amtran_bus):

Thanks.

OpenStudy (amtran_bus):

Thats understandable!

OpenStudy (arabpride):

You're Welcome ANytime!! Understandable in a sarcastic way?? XD

OpenStudy (amtran_bus):

No no I mean the way you did It I can see why I was wrong!

OpenStudy (arabpride):

oh lol!! NP hoped i helped *^.^*